OpenStudy (anonymous):

\[\int\limits_{0}^{2}\left(\begin{matrix}3x \\ x ^{2}+4\end{matrix}\right)\]

OpenStudy (mimi_x3):

\[\int\limits\frac{3x}{x^{2}-4} dx\] Not a matrices?

OpenStudy (mimi_x3):

\[\int\limits\frac{3x}{x^{2}+4} dx\]

OpenStudy (anonymous):

yup. :) but it is a definite integration. :)

OpenStudy (mimi_x3):

let \(u=x^2+4\) => \[dx= \frac{du}{2x} \]

OpenStudy (anonymous):

answer? :)

OpenStudy (anonymous):

answer @Mimi_x3 ? :)

OpenStudy (mimi_x3):

\[\int\limits\frac{3x}{u} * \frac{du}{2x} => \int\limits\frac{3\cancel{x}}{u} * \frac{du}{2\cancel{x}} => \frac{3}{2} \int\limits\frac{1}{u} du\]

OpenStudy (shubhamsrg):

you may take 3 out then take 1/2 common such that 1/2 is out and inside there is 2x now x^2 + 4 =t so 2xdx = dt => 3/2 (dt/t) try now..

OpenStudy (shubhamsrg):

just what @Mimi_x3 says :-)

OpenStudy (mimi_x3):

Integrate that then don't forget to sub back in the \(u\). (:

OpenStudy (anonymous):

ok thanks @Mimi_x3 :)